# Neural Kinematic Bases for Fluids

### v1

@@
Abstract
We propose mesh-free fluid simulations that exploit a kinematic neural basis for velocity fields represented by an MLP. We design a set of losses that ensures that these neural bases satisfy fundamental physical properties such as orthogonality, divergence-free, boundary alignment, and smoothness. Our neural bases can then be used to fit an input sketch of a flow, which will inherit the same fundamental properties from the bases. We then can animate such flow in real-time using standard time integrators. Our neural bases can accommodate different domains and naturally extend to three dimensions.
Figure 1: We use neural kinematic bases (top) to fit an input sketch of a flow (middle, left) which we then simulate in real-time with standard semi-implicit advection. We extend the same construction to three dimensions where we can simulate fluids in real time (bottom).
1 Introduction
Animating 2D or 3D shapes from sparse data is a complex problem at the core of computer graphics which started with pioneering work such as that of Badler and Morris (1982). The main motivation is character animation, where the complex animation and deformation are driven by a skeleton (sparse data) Lewis et al. (2000). The general problem can be abstracted as computing a sparse representation (or basis) that allows interactive animation and shape deformation. The main challenge is to ensure that any generated deformation looks physically realistic, which can be enforced by carefully designing the bases. With the advent of machine learning in character animation, the effort moved from creating bases to designing networks for such bases Li et al. (2021); Bertiche et al. (2021); Kavan et al. (2024).
In parallel, machine learning enabled replacing complicated and costly accurate physical simulations by cheap networks Du (2023), in particular for fluid simulations (Section 2). However, these methods share the same “problems” as physical simulations: the input is the state of the system (initial and boundary condition), and the output is an animation. This setup lacks the fine-grained control present in animation pipelines and the ability to control it from sparse data.
@@ 2 Related Work
von Tycowicz et al. (2013) accelerated the construction of reduced dynamical systems by approximating reduced forces. Our work also incorporates approximations, using smoothed boundary indicators and domain masks. Additionally, De Witt et al. (2012) represented fluids as linear combinations of eigenfunctions of the vector Laplacian, performing time integration via Galerkin discretization of the Navier-Stokes equations. While this method supported immersed rigid bodies by projecting out velocity components corresponding to boundary flux, Cui et al. (2018) extended the approach to handle Dirichlet and Neumann boundary conditions. They further improved scalability, following Jones et al. (2016), by utilizing sine and cosine transforms to reduce storage demands for eigenfunctions. Inspired by Laplacian eigenfunction methods, Mercier and Nowrouzezahrai (2020) and Wicke et al. (2009) developed reduced basis functions on regular tiles, enforcing consistency constraints between adjacent tiles. Our approach adopts a basis-function perspective but employs neural representations instead of eigenfunctions.
@@
Romero et al. (2021) augmented linear handle-based subspace formulations with nonlinear, learning-based corrections to decouple internal and external contact-driven effects. In our work, contact handling is achieved by conditioning neural basis functions on contextual information. Likewise, Aigerman et al. (2022) used neural networks to predict piecewise linear mappings of arbitrary meshes, incorporating smoothing techniques to handle discontinuities. We adopt a similar gradient-smoothing strategy to address these challenges.
@@ 3 Method
$\int_{\Omega}\text{div}(\varphi_{k})=0,\qquad\forall k=1\dots b.$
@@
$\int_{\partial\Omega}\langle n,\varphi_{k}\rangle=0,\qquad\forall k=1\dots b,$
@@
Figure 2: Example of neural basis functions generated by our MLP network. We clearly see that they are parallel to the boundary, non-zero, and smooth.
@@
Figure 3: Overview of the pipeline of our method. We start with a sketch and our pre-trained MLP that generates fluid bases. We fit the bases to the input sketch which we then advect to generate an animation.
Our goal is to learn a set of neural basis functions $\varphi_{i}$ dependent on both the position of the circles (which define the domain) and the evaluation point, such that we can use (1) as the kinematic neural basis for fluid simulation. By doing so, we can quickly evaluate specific bases (as the evaluation is done at inference time) for a given domain and thus generate fluid animation in real-time. Our design pipeline (Figure 3) starts with an input sketch of the domain $\Omega$ with a set of curves to guide the flow which we fit to our neural bases (Section 3.3). To obtain the neural bases, we define a set of losses (Section 3.1) that produces bases that satisfy fluid invariants, which we train on different domains (Section 3.2). Finally, we advect the initial fluid (Section 3.4) to generate an animation of the fluid.
3.1 Losses
We will proceed by defining a set of losses that measure the physical appropriateness of our kinematic neural basis. Our basis neural fields $\varphi_{i}$ will be computed by minimizing these losses in aggregate over a collection of $n$ randomly sampled points $p_{i}$. This sampling can be seen as using Monte Carlo integration. The main advantage of our approach is that we purely rely on fundamental physical properties (e.g., divergence-free or slip boundary conditions) and do not require any ground-truth data; we only require the bases to generalize over the position and size of the circles.
@@ 3.1 Losses
Figure 4: Boundary indicator function $w_{b}$ (left) and domain mask $w$ (right).
@@
$d_{b}(p)=\bigg{(}\frac{2\lvert d(p)\rvert^{3}}{\varepsilon^{3}}-\frac{3\lvert d(p)\rvert^{2}}{\varepsilon^{2}}+1\bigg{)}$
$w_{b}(p)=\begin{cases}0,&\lvert d(p)\rvert>\epsilon\text{ or }d_{b}(p)<0\\ d_{b}(p)^{4}&\text{otherwise}\end{cases}$
where $d(p)$ is the distance between $p$ and the boundary of $\Omega$. Then, we define the mask $w$ (Figure 4, right) by setting all values of $w_{b}$ out of the domain to zero and the values in the domain to 1. This function is one inside the domain and drops to zero for any point farther than $\varepsilon$ to the boundary. Note that the function is non-zero on a small region outside the domain, and therefore, we approximate our integral on the $[-\varepsilon,1+\varepsilon]^{2}$ domain. To ensure that our losses are independent from the number of points and their position, we normalize them by
$S=Wb,\qquad\text{with}\qquad W=\sum_{i=1}^{m}w(p_{i}).$
We can now encode the divergence invariant as the loss
$\mathcal{L}_{\text{div}}(\theta)={\sum_{k=1}^{b}\sum_{i=1}^{n}\text{div}(\varphi_{k}(p_{i}))^{2}w(p_{i})}/S.$ (2)
@@
Next, we formulate a loss for the slip boundary conditions
$\mathcal{L}_{\text{bc}}(\theta)={\sum_{k=1}^{b}\sum_{i=1}^{n}\text{cossim}(\varphi_{k}(p_{i}),n(p_{i}))^{2}w_{b}(p_{i})}/S_{b}$ (3)
@@
$\mathcal{L}_{\text{orth}}(\theta)=\sum_{i=1}^{n}\sum_{k=1}^{b}\sum_{l=k+1}^{b}\langle\varphi_{k}(p_{i}),\varphi_{l}(p_{i})\rangle w(p_{i})/S_{o}.$ (4)
@@
$\mathcal{L}_{\text{len}}(\theta)={\sum_{k=1}^{b}\bigg{(}\sum_{i=1}^{n}\|\varphi_{k}(\theta,p_{i})\|w(p_{i})/W-c\bigg{)}^{2}}/b,$ (5)
@@
$\mathcal{L}_{\text{small}}(\theta)=\sum_{k=1}^{b}\sum_{i=1}^{n}\text{ReLU}(\delta-\|\varphi_{k}(p_{i})\|)w(p_{i})/S.$ (6)
@@
$\mathcal{L}_{\text{smooth}}(\theta)={\sum_{k=1}^{b}\sum_{i=1}^{n}\|J_{\varphi_{k}}(p_{i})\|_{F}^{2}w(p_{i})}/S,$ (7)
@@
3.2 Training
We limit the training to a domain represented by three circles and compute the distance $d(p)$ using a soft minimum to the 3 circles. Our neural fields are parameterized by MLPs that have 8 fully connected layers, which use the leaky ReLU activation function (except ELU for the last layer as we want to produce negative numbers) and have 256 channels per layer (Figure 5). The input vector is the concatenation of the position of one sample point $p_{i}$ and 3 circle parameters. The output vector consists of the vector for the evaluation of the $b$ neural bases.
Figure 5: Overview of our MLP architecture: it takes as input a point $p$ and the set of circles’ parameters $\rho$ and produces our neural bases $\varphi_{i},i=1,\dots,b$.
The weights for terms in the fluid loss are $w_{\text{drch}}:w_{\text{div}}:w_{\text{orth}}:w_{\text{bc}}:w_{\text{len}}:w_{\text{small}}=0.01:5:100:50:100:100$. We choose the threshold for small base penalty $\delta$ as 0.05 and the average length for length penalty as $c=0.37$.
The MLP uses Kaiming initialization for its parameters. We train the MLP using the Adam optimizer with a learning rate initialized at 0.001. To dynamically adjust the learning rate, we employed an ExponentialLR scheduler, applying a decay factor of 0.96, resulting in a learning rate scaled by $LR\times 0.96^{t}$ per epoch, where $t$ is the epoch.
We train on 1000 different geometric samples, each of which is a set of randomly generated 3 circles with centers in the $[0.25,0.75]^{2}$ square and radius in the $[0.1,0.3]$ interval. They can be (partially) overlapping or separate. For each sample, we randomly sample $n=10^{6}$ points over the domain. The training was conducted for a total of 10 epochs with a batch size of $10^{4}$.
We train the boundary loss in two stages: first, we use a power of 4 in the cosine similarity, then lower it to 2. This approach leads the MLP to first remove outliers and have a correct alignment close to the boundary, Figure 6 left (higher power leads to a narrower band). By then lowering it, the MLP can relax the condition and propagate the aligned to a larger (and smoother region), Figure 6, right.
Figure 6: Visualization of three bases at the end of the first stage when the power is four, top. After lowering the power to two, the bases align well with the boundary, bottom.
3.3 Fitting to a Sketch
@@
Figure 7: Example of the different interactions possible to design a fluid flow. From the initial design to different types of editing.
@@
3.4 Advection
We then use $\alpha^{0}$ to compute the initial velocity $v^{0}$, which we advect using a semi-implicit integrator Stam (1999, 2023). For every time step $t$, for every point, we compute the origin position
@@ 3.4 Advection
$\alpha^{t+1}=\operatorname*{arg\,min}_{\alpha}\sum_{i}\left\|\sum_{k=1}^{b}\varphi_{k}(p_{i})\alpha_{k}-\bar{v}^{t+1}(p_{i})\right\|^{2},$
@@
4 Results
@@
4.1 Generalization
We evaluate our neural bases on a test dataset comprised of 100 random unseen circles with centers and radii sampled from the same distribution (figures 8-10, purple, show the value of the different losses during training). At the end of the training, our neural bases have similar losses: the length losses and orthogonality are practically the same, while the divergence and boundary loss are just around 15% larger.
4.2 Convergence
Overall, our fluid loss steadily decreased over the training process. Note that we use blue vertical lines to indicate when the second stage of our training starts. After an initial phase, both length losses and orthogonality become stationary, indicating that the bases reached the desired length, are individually not zero, and are different (Figure 8). It is interesting to note that subsequent epochs do not change their value, indicating that the training mostly rotates the vectors. The orthogonality loss and the smoothness loss (which depend on the gradient of the MLP and impose first-order conditions) require more steps (400 thousand) to reach a stationary value, suggesting that our bases maintain the fluid properties and are smooth in most of our training process (Figure 9). Both losses are affected by the two-stage training, but they quickly reach a new stable flat state. We note that, as expected, the smoothness loss increases in the beginning as the bases become larger. Additionally, the smoothness loss is the only loss that can not reach a minimum of zero, as it would lead to constant bases. The boundary condition loss is the only one affected by the two-stage training: after the first 10 epochs, we decrease the power of the cosine similarity (Figure 10). We clearly see the “jump” in the loss when we switch. After the initial increase, the loss quickly decreases and reaches similar values.
@@ 4.2 Convergence
Figure 8: Convergence of the length-based losses. After the first 100 thousand steps, both losses are converged.
@@
Figure 9: Convergence of the “gradient” based losses. These losses impose conditions on the gradient and thus require longer to converge.
@@
Figure 10: Convergence of the boundary loss. When decreasing the power (blue line), the loss increases to then quickly drops again.
4.3 Ablation study

### v2

@@
Abstract.
We propose mesh-free fluid simulations that exploit a kinematic neural basis for velocity fields represented by an MLP. We design a set of losses that ensures that these neural bases approximate fundamental physical properties such as orthogonality, divergence-free, boundary alignment, and smoothness. Our neural bases can then be used to fit an input sketch of a flow, which will inherit the same fundamental properties from the bases. We then can animate such flow in real-time using standard time integrators. Our neural bases can accommodate different domains, moving boundaries, and naturally extend to three dimensions.
Real-Time Fluid Animation, Reduced Order Method, Physics Informed Neural Networks
journal: TOGccs: Computing methodologies Physical simulationccs: Computing methodologies Neural networksccs: Computing methodologies Interactive simulationccs: Computing methodologies Simulation by animation
Figure 1. We use neural kinematic bases (top) to fit a changing domain which we then simulate in real-time with standard semi-implicit advection (middle). The same construction also applied to two dimensions where we can simulate fluids in real time (bottom).
1. Introduction
Animating 2D or 3D shapes from sparse data is a complex problem at the core of computer graphics which started with pioneering work such as that of Badler and Morris (1982). The main motivation is character animation, where the complex animation and deformation are driven by a skeleton (sparse data) (Lewis et al., 2000). The general problem can be abstracted as computing a sparse representation (or basis) that allows interactive animation and shape deformation. The main challenge is to ensure that any generated deformation looks physically realistic, which can be enforced by carefully designing the bases. With the advent of machine learning in character animation, the effort moved from creating bases to designing networks for such bases (Li et al., 2021; Bertiche et al., 2021; Kavan et al., 2024).
In parallel, machine learning enabled replacing complicated and costly accurate physical simulations by cheap networks (Du, 2023), in particular for fluid simulations (Section 2). However, these methods share the same “problems” as physical simulations: the input is the state of the system (initial and boundary condition), and the output is an animation. This setup lacks the fine-grained control present in animation pipelines and the ability to control it from sparse data.
@@ 2. Related Work
von Tycowicz et al. (2013) accelerated the construction of reduced dynamical systems by approximating reduced forces; Umetani and Bickel (2018) uses data-drive approach to quickly predict fluid behavior around 3D obstacles. Our work also incorporates approximations, using smoothed boundary indicators and domain masks. Additionally, De Witt et al. (2012) represented fluids as linear combinations of eigenfunctions of the vector Laplacian, performing time integration via Galerkin discretization of the Navier-Stokes equations. While this method supported immersed rigid bodies by projecting out velocity components corresponding to boundary flux, Cui et al. (2018) extended the approach to handle Dirichlet and Neumann boundary conditions. They further improved scalability, following Jones et al. (2016), by utilizing sine and cosine transforms to reduce storage demands for eigenfunctions. Inspired by Laplacian eigenfunction methods, Mercier and Nowrouzezahrai (2020) and Wicke et al. (2009) developed reduced basis functions on regular tiles, enforcing consistency constraints between adjacent tiles. Our approach adopts a basis-function perspective but employs neural representations instead of eigenfunctions. Comparing to Cui et al. (2018), which is limited to generating bases in a single rectangular domain and handles obstacle boundaries via an explicit post-processing step, our neural network is trained to generalize across arbitrary domains and evaluates bases on-the-fly during simulation. While Cui et al. (2018) require at least 5.5 hours of precomputation per domain, our model is trained at a one-time cost, after which it supports unlimited simulations without domain-specific precomputation. By treating both obstacles and rectangular boundaries consistently during basis generation, our method naturally satisfies no-slip boundary conditions (Appendix A).
@@
Romero et al. (2021) augmented linear handle-based subspace formulations with nonlinear, learning-based corrections to decouple internal and external contact-driven effects. In our work, contact handling is achieved by conditioning neural basis functions on contextual information. Likewise, Aigerman et al. (2022) used neural networks to predict piecewise linear mappings of arbitrary meshes, incorporating smoothing techniques to handle discontinuities. We adopt a similar gradient-smoothing strategy to address these challenges. Similar to Sharp et al. (2023), which presents a reduced-order simulation method using unsupervised learning for neural network training, we design a set of physics-informed losses, inspired by Karniadakis et al. (2021), to enable unsupervised training of our neural network.
@@ 3. Method
$\text{div}(\varphi_{k})=0,\qquad\forall k=1\dots b.$
@@
$\langle n,\varphi_{k}\rangle=0,\qquad\forall k=1\dots b,$
@@
Figure 2. Example of neural basis functions generated by our MLP network. We clearly see that they are parallel to the boundary, non-zero, and smooth.
@@
Figure 3. Overview of the pipeline of our method. We start with a sketch and our pre-trained MLP that generates fluid bases. We fit the bases to the input sketch which we then advect to generate an animation.
Our goal is to learn a set of neural basis functions $\varphi_{i}$ dependent on both the domain (which we encode with a set of circles) and the evaluation point, such that we can use (1) as the kinematic neural basis for fluid simulation. By doing so, we can quickly evaluate specific bases (as the evaluation is done at inference time) for a given domain and thus generate fluid animation in real-time. Our design pipeline (Figure 3) starts with an input sketch of the domain $\Omega$ with a set of curves to guide the flow which we fit to our neural bases (Section 3.4). To obtain the neural bases, we define a set of losses (Section 3.2) that produces bases that approximate fluid invariants in average, which we train (unsupervised) on different domains (Section 3.3). Finally, we advect the initial fluid (Section 3.5) to generate an animation of the fluid.
3.1. Domain
We encode the domain with an unit square with rounded corners with radius 0.2 and an obstacle blob. To keep the representation simple, we use ten implicit circles (or spheres in 3d)
$f_{i}(p)=\|p-c_{i}\|^{2}-r_{i}^{2}\quad\text{with}i=1,\dots,10,$
where $p$ is a point on the plane (or space), $c_{i}$ is the circle’s center, and $r_{i}$ is its radius. We combine them by taking inspiration from metaballs or blobs (Blinn, 1982) to generate the implicit function
$f(p)=\ln\big(\sum_{i=1}^{10}e^{-kf_{i}(p)}\big)/k,$
where $k$ controls the smoothness (in our results, we always use $k=30$).
$d(p)=\bigg(\frac{2\lvert f(p)\rvert^{3}}{\varepsilon^{3}}-\frac{3\lvert f(p)\rvert^{2}}{\varepsilon^{2}}+1\bigg)$
$w_{b}(p)=\begin{cases}0,&\lvert f(p)\rvert>\epsilon\text{ or }d(p)<0\\ d(p)^{4}&\text{otherwise}.\end{cases}$
Then, we define the mask $w$ (Figure 4, right) by setting all values of $w_{b}$ out of the domain to zero and the values in the domain to 1. This function is one inside the domain and drops to zero for any point farther than $\varepsilon$ from the boundary. Note that the function is non-zero on a small region outside the domain, and therefore, we approximate our integral on the $\varepsilon,1+\varepsilon]^{2}$ domain. To ensure that our losses are independent from the number of points and their position, we normalize them by
$S=Wb,\qquad\text{with}\qquad W=\sum_{i=1}^{m}w(p_{i}).$
3.2. Losses
We will proceed by defining a set of losses that measure the physical appropriateness of our kinematic neural basis. Our basis neural fields $\varphi_{i}$ will be computed by minimizing these losses in aggregate over a collection of $n$ randomly sampled points $p_{i}$. This sampling can be seen as using Monte Carlo integration. The main advantage of our approach is that we purely rely on fundamental physical properties (e.g., divergence-free or slip boundary conditions) and do not require any ground-truth data; we only require the bases to generalize over the domain.
@@ 3.2. Losses
Figure 4. Boundary indicator function $w_{b}$ (left) and domain mask $w$ (right).
@@
We can now encode the divergence invariant as an average, using the loss
(2) $\mathcal{L}_{\text{div}}(\theta)={\sum_{k=1}^{b}\sum_{i=1}^{n}\text{div}(\varphi_{k}(p_{i}))^{2}w(p_{i})}/S.$
@@
Next, we formulate a loss for the average slip boundary conditions
(3) $\mathcal{L}_{\text{bc}}(\theta)={\sum_{k=1}^{b}\sum_{i=1}^{n}\text{cossim}(\varphi_{k}(p_{i}),n(p_{i}))^{2}w_{b}(p_{i})}/S_{b}$
@@
(4) $\mathcal{L}_{\text{orth}}(\theta)=\sum_{i=1}^{n}\sum_{k=1}^{b}\sum_{l=k+1}^{b}\langle\varphi_{k}(p_{i}),\varphi_{l}(p_{i})\rangle w(p_{i})/S_{o}.$
@@
(5) $\mathcal{L}_{\text{len}}(\theta)={\sum_{k=1}^{b}\bigg(\sum_{i=1}^{n}\|\varphi_{k}(\theta,p_{i})\|w(p_{i})/W-c\bigg)^{2}}/b,$
@@
(6) $\mathcal{L}_{\text{small}}(\theta)=\sum_{k=1}^{b}\sum_{i=1}^{n}\text{ReLU}(\delta-\|\varphi_{k}(p_{i})\|)w(p_{i})/S.$
@@
(7) $\mathcal{L}_{\text{smooth}}(\theta)={\sum_{k=1}^{b}\sum_{i=1}^{n}\|J_{\varphi_{k}}(p_{i})\|_{F}^{2}w(p_{i})}/S,$
@@
3.3. Training
We train on a domain represented by ten circles. Our neural fields are parameterized by MLPs that have 8 fully connected layers, which use the leaky ReLU activation function (except ELU for the last layer as we want to produce negative numbers) and have 256 channels per layer (Figure 5). The input vector is the concatenation of the position of one sample point $p_{i}$ and 10 circle parameters. The output vector consists of the vector for the evaluation of the $b$ neural bases.
Figure 5. Overview of our MLP architecture: it takes as input a point $p$ and the set of circles’ parameters $\rho$ and produces our neural bases $\varphi_{i},i=1,\dots,b$.
We assign the following weights to each term in the fluid loss function: $(w_{\text{drch}},w_{\text{div}},w_{\text{orth}},w_{\text{bc}},w_{\text{len}},w_{\text{small}})={(0.01,5,100,30,100,100)}$. We choose the threshold for small base penalty $\delta$ as 0.05 and the average length for length penalty as $c=0.37$.
The MLP uses Kaiming initialization for its parameters. We train the MLP using the Adam optimizer with a learning rate initialized at 0.0005. To dynamically adjust the learning rate, we employed an ExponentialLR scheduler, applying a decay factor of 0.96, resulting in a learning rate scaled by $LR\times 0.96^{t}$ per epoch, where $t$ is the epoch.
We train on 1000 different geometric samples, containing randomly generated circles for 2D and spheres for 3D with centers in the $0.225,0.675]^{d}$ square or cube and radius in the $[0.03,0.09]$ interval, randomly generated shapes with two and three components, and, in 2D, shapes representing the letters of the english alphabet. They can be (partially) overlapping or separate. For each sample, we randomly sample $n=10^{6}$ points over the domain. The training was conducted for a total of 10 epochs with a batch size of $10^{4}$.
The implementation is based on PyTorch, with a training time of 23.6 hours for the 3D model and 22.9 hours for the 2D model, on a single NVIDIA GeForce RTX 3090 GPU.
3.4. Fitting to a Sketch
@@
Figure 6. Example of the different interactions possible to design a fluid flow. From the initial design to different types of editing.
@@
3.5. Advection
We then use $\alpha^{0}$ to compute the initial velocity $v^{0}$, which we advect using a semi-implicit integrator (Stam, 1999, 2023). For every time step $t$, for every point, we compute the origin position
@@ 3.5. Advection
(8) $\alpha^{t+1}=\operatorname*{arg\,min}_{\alpha}\sum_{i}\left\|\sum_{k=1}^{b}\varphi_{k}(p_{i})\alpha_{k}-\bar{v}^{t+1}(p_{i})\right\|^{2},$
@@
Moving Domain
Since our neural bases can be evaluated on arbitrary domains, they can naturally capture fluid flows with moving boundaries. In the time integrator, after computing the new velocity, we instead of projecting $\bar{v}^{t+1}(p)$ on the same basis (8), we project it on new set of bases defined on a different domain.
4. Results
@@
4.1. Generalization
We evaluate our neural bases on a test dataset comprised of 100 random unseen circles/spheres with centers and radii sampled from the same distribution (Figure 7, orange dashed, show the value of the different losses during training). At the end of the training, our neural bases have similar losses: the length losses and orthogonality are practically the same, while the boundary loss is just around 20% larger.
@@ 4.1. Generalization
2D
3D
Figure 7. Convergence of the losses over the number of epochs across the training (blue) and test (orange) set.
Divergence
Boundary alignment
2D
3D
Figure 8. Density plot of divergence loss and boundary loss on the training (blue) and test (red) sets. The $x$-axis shows loss values, and the $y$-axis indicates the probability density.
4.2. Convergence
Overall, our fluid loss steadily decreased over the training process. In 2D and 3D, after the initial few epochs the small and orthogonality loss become stationary, indicating that the bases are individually not zero, and are different (Figure 7). In 2D the bases also quickly reach the target length, while in 3D it slowly converges over the whole training process. It is interesting to note that in 2D subsequent epochs do not change their value, indicating that the training mostly rotates the vectors. The divergence and boundary loss require more epochs (10-15) to reach a stationary value, since these losses are the only once affected by the shape of the domain (Figure 7). We note that, as expected, the smoothness loss increases in the beginning as the bases become larger. To evaluate the plausibility of our bases, we measure the distribution of the divergence and boundary alignment on training (blue) and test (red) set (Figure 8). The divergence is consistently small (but not zero) across the two sets (in average 0.5), while the boundary distribution is slightly larger for the test set.
4.3. Ablation study